OpenStudy (anonymous):

Simplify by removing factors of one. (q^2-16)/(q^2-8q+16) Is the answer 8q?

OpenStudy (anonymous):

\[\frac{\left(q^2-16\right)}{\left(q^2-8q+16\right)}=\frac{(q-4) (q+4)}{(q-4)^2}=\frac{q+4}{q-4} \]